I need help solving this:
Find 3 consecutive integers 4 times the 3rd decreased by 1/2 second = 4

Thank you so much in advance.
Dan

Answer by stanbon(75887)

Find 3 consecutive integers 4 times the 3rd decreased by 1/2 second = 4
---
1st: x-2
2nd: x-1
3rd: x
4th: x+1
------------------------------
Equation:
4x - (1/2)(x-1) = 4
---
8x - x + 1 = 8
7x = 7
x = 1
----
1st: x-2 = -1
2nd: x-1 = 0
3rd: x = 1
4th: x = 2
=================
Cheers,
Stan H.